# Bayesian Inference Under Differential Privacy: Prior Selection Considerations with Application to Univariate Gaussian Data and Regression

### v1

@@ Abstract
We describe Bayesian inference for the mean and variance of bounded data protected by differential privacy and modeled as Gaussian. Using this setting, we demonstrate that analysts can and should take the constraints imposed by the bounds into account when specifying prior distributions. Additionally, we provide theoretical and empirical results regarding what classes of default priors produce valid inference for a differentially private release in settings where substantial prior information is not available. We discuss how these results can be applied to Bayesian inference for regression with differentially private data.
@@ 1 Introduction
Differential privacy (DP) [12] is the current gold standard for protecting individual privacy when performing data releases involving confidential data. A common method of ensuring a release satisfies DP is to add calibrated random noise to sufficient statistics of the data. Once the noisy statistics are released, analysts in theory can use them to perform statistical tasks such as the estimation of unknown parameters of interest, prediction of new data values, and quantification of the uncertainty around these quantities. Performing these tasks in a principled way, however, can be nontrivial and is an area of active research.
@@
Bayesian methods are natural for statistical inference after a DP release. By sampling from posterior distributions, analysts can perform estimation and prediction tasks automatically with “built-in” uncertainty quantification. As examples, Bernstein and Sheldon (2018) [4] and Ju et al. (2022) [19] use Gibbs sampling to estimate differentially private posterior distributions, the former for exponential families and the latter for likelihoods satisfying a record additivity condition. Gong (2022) [16] uses approximate Bayesian computation for settings where a perturbation mechanism is applied. Other works investigate Bayesian estimation of subset proportions [28], linear regression coefficients [5], and Poisson factorization [36]. These methods all employ variations of the following scheme: repeatedly impute plausible values of the nonprivate statistics based on their private counterparts, and obtain samples of the parameters of interest via a traditional nonprivate Bayesian analysis.
@@
DP is related to Bayesian methods more generally. For example, releasing a sample from a posterior distribution can be used to achieve DP [9, 15, 18, 31, 40, 42, 43]. Another line of work examines Bayesian semantics for DP and their relationship to disclosure risk assessment [12, 21, 22, 23, 25, 26, 29]. We also note that researchers have developed non-Bayesian methods for inference under DP, including interval estimates [2, 10, 11, 13, 20, 37] and significance tests [1, 2, 3, 7, 24, 35].
@@
Existing methods for Bayesian inference under DP generally presume that the analyst sets an informative prior distribution. In practice, however, the analyst may not have sufficient prior knowledge to set an accurate, informative prior. Additionally, methods in the literature typically rely on prior distributions with unbounded support. Often, however, the confidential data are assumed to be constrained in some way; for example, values lie within an interval $[a,b]$ [5, 11, 37, 39, 41]. Arguably, samples from the posterior distribution should reflect this constraint as well. Yet, existing methods for Bayesian inference under DP can produce estimates or predictions outside the feasible range. This may lead analysts to resort to ad hoc fixes such as clipping estimates or predictions to be at the bounds of the interval, which can lead to undesirable and unpredictable behavior.
@@
We address issues resulting from ignoring constraints in Bayesian inference under DP and the use of default prior distributions under DP. Our primary contributions include:
@@
    We illustrate the incorporation of constraints into a differentially private, Bayesian inference procedure and demonstrate their effect on estimation and prediction tasks.
@@
    We present a first-of-its-kind examination of default prior distributions analysts might consider in lieu of highly informative proper priors, and present theoretical results demonstrating the need for caution when certain weakly informative priors are used in a DP analysis.
Throughout, we use the univariate Gaussian setting as a motivating case study. This setting is among the most important and commonly encountered in statistics but is relatively understudied in the DP literature, with only a few works [10, 11, 13, 20] providing frequentist methods for inference with uncertainty quantification. General Bayesian DP inference methods such as those in [4, 16, 19] can be applied to this setting but require approximations or substantial computation time. As part of our contributions, we propose a Gibbs sampling procedure for the univariate Gaussian setting without these drawbacks.
2 Background and Case Study
In this section, we briefly review the definition of DP and several of its properties. We then describe the DP univariate Gaussian setting and present our novel Gibbs sampling scheme.
@@ 2.1 Differential Privacy
As described by [12], a release mechanism ${\mathcal{M}}$ satisfies $\varepsilon$-DP if for all data sets $x,x^{\prime}$ that differ in only one row and all $S\subseteq Range({\mathcal{M}})$,
@@
DP has many desirable properties [12]. Here, we make use of post-processing. For any data set $x$, if ${\mathcal{M}}(x)$ satisfies $\varepsilon$-DP and $g$ is not a function of $x$, then $g({\mathcal{M}}(x))$ satisfies $\varepsilon$-DP. We also make use of composition. For any data set $x$, if $m_{1}={\mathcal{M}}_{1}(x)$ satisfies $\varepsilon_{1}$-DP and $m_{2}={\mathcal{M}}_{2}(x,m_{1})$ satisfies $\varepsilon_{2}$-DP, then ${\mathcal{M}}(x)=(m_{1},m_{2})$ satisfies $(\varepsilon_{1}+\varepsilon_{2})$-DP.
@@ 2.2 The Univariate Gaussian Setting
Let $Y_{1},\ldots,Y_{n}$ be the confidential, scalar data values for $n$ individuals. We assume that all $Y_{i}$ are contained in some publicly known interval $[a,b]$. The analyst models the data as independent and identically distributed with $Y_{i}\sim{\mathcal{N}}(\mu,\sigma^{2})$ for all $i$. Gaussian models are reasonable and commonly employed when the observed data are symmetric with a range not near the bounds. For example, quantities like proportions and standardized test scores have known ranges. Other quantities like commute times and stock market returns have known lower bounds and effective upper bounds.
The sufficient statistics are $\bar{Y}=\sum_{i=1}^{n}Y_{i}/n$ and $S^{2}=\sum_{i=1}^{n}(Y_{i}-\bar{Y})^{2}/(n-1)$. As shown in [11], $\bar{Y}$ and $S^{2}$ have sensitivities $(b-a)/n$ and $(b-a)^{2}/n$, respectively. To achieve $\varepsilon$-DP, the data curator can release $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},(b-a)/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},(b-a)^{2}/(\varepsilon_{2}n))$, where $\varepsilon_{1}+\varepsilon_{2}=\varepsilon$. We make the simplifying assumption that $a=0$ and $b=1$. We can do so without loss of generality, as described in Theorem 1, which is proved in Appendix A.
@@
For now, we suppose the analyst wishes to incorporate prior beliefs into their analysis via an informative prior distribution. A conjugate choice for the prior in the public setting is
$\sigma^{2}\sim\textsf{IG}\left(\nu_{0}/2,\nu_{0}\sigma_{0}^{2}/2\right),\qquad\mu\mid\sigma^{2}\sim{\mathcal{N}}\left(\mu_{0},\sigma^{2}/\kappa_{0}\right)$ (2)
with analyst-specified hyperparameters $(\nu_{0},\kappa_{0},\mu_{0},\sigma_{0}^{2}).$
We now present a Gibbs sampler to estimate this model, building on the ideas in [4, 5]. Our innovation is with respect to the full conditional for $S^{2}$, for which we derive an exact form. As long as $\varepsilon_{2}<2$, the full conditional for $S^{2}$ is a distribution we call a truncated gamma mixture (TGM). [^1] For the distribution of $S^{2}$ to be a TGM, we require $\sigma^{2}<(n-1)/(2n\varepsilon_{2})$. We show in Corollary 1 that in this setting we can guarantee $\sigma^{2}\leq 1/4$, which implies $\sigma^{2}<(n-1)/(2n\varepsilon_{2})$ automatically when $\varepsilon_{2}<2$. Within the Gibbs sampler, we reject and resample any draws for which $\sigma^{2}\geq(n-1)/(2n\varepsilon_{2})$.
@@
A random variable $X\sim\textsf{TGM}(\alpha,\beta,\lambda,\tau)$ for $\alpha>0$, $\beta>\lambda\geq 0$, and $\tau\in{\mathbb{R}}$ if it has probability density function
$p(x)=\begin{cases}\frac{(\beta+\lambda)^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-(\beta+\lambda)x},&\mbox{if }\tau\leq 0;\\ \pi_{1}\frac{(\beta-\lambda)^{\alpha}}{\gamma(\alpha,(\beta-\lambda)\tau)}x^{\alpha-1}e^{-(\beta-\lambda)x},&\mbox{if }\tau>0\mbox{ and }x\leq\tau;\\ \pi_{2}\frac{(\beta+\lambda)^{\alpha}}{\Gamma(\alpha,(\beta+\lambda)\tau)}x^{\alpha-1}e^{-(\beta+\lambda)x}&\mbox{if }\tau>0\mbox{ and }x>\tau,\end{cases}$ (3)
@@
$\pi_{1}=\frac{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}}{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}+e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}},\qquad\pi_{2}=\frac{e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}}{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}+e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}}.$ (4)
@@
Here, $\gamma(\alpha,x)$ is the lower incomplete gamma function and $\Gamma(\alpha,x)$ is the upper incomplete gamma function. We provide an algorithm for sampling from the TGM in Appendix B.
@@
Derivations for all full conditionals for the Gibbs sampler are provided in Appendix C. The full conditionals for $\mu$ and $\sigma^{2}$ are identical to those of the nonprivate setting. To obtain the full conditional for $\bar{Y}$, we follow the strategy of [4] based on the Bayesian LASSO [34] and use the fact that $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ is equal in distribution to $\bar{Y}^{*}\sim{\mathcal{N}}(\bar{Y},\omega^{2})$ for $\omega^{2}\sim\textsf{Exp}(\varepsilon_{1}^{2}n^{2}/2)$. The full conditional for $\bar{Y}^{*}$ is Gaussian, the full conditional for $1/\omega^{2}$ is inverse-Gaussian, and the full conditional for $S^{2}$ is a TGM. The parameters of the full conditionals are provided in Appendix D.
@@
A major advantage of this sampler is its runtime. If $T$ is the number of Gibbs iterations, then the computational complexity of this sampler is $\mathcal{O}(T)$; the runtime does not scale with the sample size, $n$. The Gibbs sampler of [4] for exponential families applied to the univariate Gaussian setting also has computational complexity $\mathcal{O}(T)$. This sampler, however, requires a Gaussian approximation for $S^{2}$, which, particularly when $\sigma^{2}$ is near zero or $n$ is small, can lead to unreliable results. The sampler of [19] does not require any approximations; however, it has computational complexity $\mathcal{O}(nT)$, making it potentially infeasible for large $n$. Our proposed sampler achieves both the exact sampling of [19] and the $\mathcal{O}(T)$ computational complexity of [4].
3 Incorporating Constraints

### v2

# Bayesian Inference Under Differential Privacy With Bounded Data

@@ Abstract
We describe Bayesian inference for the parameters of Gaussian models of bounded data protected by differential privacy. Using this setting, we demonstrate that analysts can and should take constraints imposed by the bounds into account when specifying prior distributions. Additionally, we provide theoretical and empirical results regarding what classes of default priors produce valid inference for a differentially private release in settings where substantial prior information is not available. We discuss how these results can be applied to Bayesian inference for regression with differentially private data.
@@ 1 INTRODUCTION
Differential privacy (DP) (Dwork et al., 2006) is the current gold standard for protecting individual privacy when performing data releases involving confidential data. A common method of ensuring a release satisfies DP is to add calibrated random noise to sufficient statistics of the data. Once the noisy statistics are released, analysts in theory can use them to perform statistical tasks such as the estimation of unknown parameters of interest, prediction of new data values, and quantification of the uncertainty around these quantities. Performing these tasks in a principled way, however, can be nontrivial and is an area of active research.
@@
Bayesian methods are natural for statistical inference after a DP release. By sampling from posterior distributions, analysts can perform estimation and prediction tasks automatically with “built-in” uncertainty quantification. As examples, Bernstein & Sheldon (2018) and Ju et al. (2022) use Gibbs sampling to estimate differentially private posterior distributions, the former for exponential families and the latter for likelihoods satisfying a record additivity condition. Gong (2022) uses approximate Bayesian computation for settings where a perturbation mechanism is applied. Other works investigate Bayesian estimation of subset proportions (Li & Reiter, 2022), regression coefficients (Bernstein & Sheldon, 2019), and Poisson factorization (Schein et al., 2019). These methods employ variations of the following scheme: repeatedly impute plausible values of the nonprivate statistics based on their private counterparts, and obtain samples of the parameters of interest via a nonprivate Bayesian analysis.
@@
DP is related to Bayesian methods more generally. Releasing a sample from a posterior distribution can be used to achieve DP (Dimitrakakis et al., 2017; Geumlek et al., 2017; Jewson et al., 2024; Minami et al., 2016; Wang et al., 2015; Zhang & Zhang, 2023; Zhang et al., 2016). Another line of work examines Bayesian semantics for DP and their relationship to disclosure risk assessment (Dwork et al., 2006; Kasiviswanathan & Smith, 2014; Kazan & Reiter, 2023, 2024; Kifer et al., 2022; Kifer & Machanavajjhala, 2014; McClure & Reiter, 2012).
@@
Existing methods for Bayesian inference under DP generally presume that the analyst sets an informative prior distribution. In practice, however, the analyst may not have sufficient prior knowledge to set an accurate, informative prior. Additionally, methods in the literature typically rely on prior distributions with unbounded support. Often, however, the confidential data are assumed to be constrained in some way; for example, values lie within an interval $[a,b]$ (Bernstein & Sheldon, 2019; Du et al., 2020; Sheffet, 2017; Wang, 2018; Zhang et al., 2012). Arguably, samples from the posterior distribution should reflect this constraint as well. Yet, existing methods for Bayesian inference under DP can produce estimates or predictions outside the feasible range. This may lead analysts to resort to ad hoc fixes such as clipping estimates or predictions to be at the bounds of the interval, which can lead to undesirable and unpredictable behavior.
@@
We address incorporating constraints in Bayesian inference under DP, including implications for prior specification. Our primary contributions include:
@@
    We propose an improved Gibbs sampler for Gaussian models to enable posterior inference after the release of DP statistics. The sampler uses exact forms of all full conditional distributions, avoiding approximations that are typically used in other Bayesian post-processing inferential approaches.
@@
    We propose a method to incorporate constraints into a Bayesian inference procedure after a DP data release that is computationally efficient, simple, and generalizes to a range of settings. We illustrate its benefits for estimation and prediction.
  3. 3.
    We present a first-of-its-kind examination of default prior distributions analysts might consider in lieu of highly informative proper priors, and present theoretical results demonstrating the need for caution when considering weakly informative priors in a DP analysis.
Throughout, we use the univariate Gaussian setting as a motivating case study. This setting is among the most important and commonly encountered in statistics but is relatively understudied in the DP literature, with only a few works providing frequentist methods for interval estimation (D’Orazio et al., 2015; Du et al., 2020; Ferrando et al., 2022; Karwa & Vadhan, 2017) or hypothesis testing (Couch et al., 2019; Kazan et al., 2023; Peña & Barrientos, 2024).
2 BACKGROUND AND SETTING
We first briefly review DP and several of its properties. We then describe the DP univariate Gaussian setting and present our novel Gibbs sampling scheme.
@@ 2.1 Differential Privacy
As described by Dwork et al. (2006), a release mechanism ${\mathcal{M}}$ satisfies $\varepsilon$-DP if for all data sets $x,x^{\prime}$ that differ in only one row and all $S\subseteq Range({\mathcal{M}})$,
@@
DP has many desirable properties (Dwork et al., 2006). Here, we make use of post-processing. For any data set $x$, if ${\mathcal{M}}(x)$ satisfies $\varepsilon$-DP and $g$ is not a function of $x$, then $g({\mathcal{M}}(x))$ satisfies $\varepsilon$-DP. We also make use of composition. For any data set $x$, if $m_{1}={\mathcal{M}}_{1}(x)$ satisfies $\varepsilon_{1}$-DP and $m_{2}={\mathcal{M}}_{2}(x,m_{1})$ satisfies $\varepsilon_{2}$-DP, then ${\mathcal{M}}(x)=(m_{1},m_{2})$ satisfies $(\varepsilon_{1}+\varepsilon_{2})$-DP.
@@ 2.2 The Univariate Gaussian Setting
Let $Y_{1},\ldots,Y_{n}$ be confidential, scalar data values for $n$ individuals. The $Y_{i}$ are contained in some interval $[a,b]$ set by the data curator that accords with scientific understanding of the phenomenon under study and, to maintain the DP guarantee, is independent of the realized data values. As examples, test scores may be bounded between $[a,b]=[0,100]$ and IQ scores between $[a,b]=[40,250]$. This interval is made public by the data curator.
The data curator releases private versions of the sample mean and variance, i.e., $\bar{Y}=\sum_{i=1}^{n}Y_{i}/n$ and $S^{2}=\sum_{i=1}^{n}(Y_{i}-\bar{Y})^{2}/(n-1)$. As shown in Du et al. (2020), $\bar{Y}$ and $S^{2}$ have sensitivities $(b-a)/n$ and $(b-a)^{2}/n$, respectively. To achieve $\varepsilon$-DP, we presume the data curator releases $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},(b-a)/(\varepsilon_{1}n))$ and $S^{2*}\sim\textsf{Lap}(S^{2},(b-a)^{2}/(\varepsilon_{2}n))$, where $\varepsilon_{1}+\varepsilon_{2}=\varepsilon$.
We make the simplifying assumptions $a=0$ and $b=1$. We can do so without loss of generality, as described in Theorem 1, which is proved in Appendix A.
@@
A data analyst who only has access to $\bar{Y}^{*}$ and $S^{2*}$, as well as the particulars of the DP mechanism used to create them, considers these released summary statistics as noisy versions of $\bar{Y}$ and $S^{2}$, which themselves are estimates of underlying population parameters $\mu$ and $\sigma^{2}$. To make inferences for these population parameters, we presume the analyst posits a model for the underlying confidential data, namely $Y_{i}\sim{\mathcal{N}}(\mu,\sigma^{2})$ independently for all $i$. While theoretically this model has support outside the interval $[a,b]$, in practice it accurately describes data distributions when the observed values are symmetric with an empirical range not near the bounds. Indeed, data analysts frequently use normal models for bounded data; for example, normal models are typical for quantities like proportions and standardized test scores, which have known ranges. They also are used, for example, for quantities like health measurements (e.g., blood pressures, birth weights) and stock market returns, which can be bounded based on extensive historical observations and scientific considerations.
We consider an analyst who follows the Bayesian paradigm to obtain posterior inferences about $(\mu,\sigma^{2})$ given $(\bar{Y}^{*},S^{2*})$. We consider first the case of an analyst who incorporates prior beliefs into their analysis via the convenient, informative prior distribution,
$\sigma^{2}\sim\textsf{IG}\left(\nu_{0}/2,\nu_{0}\sigma_{0}^{2}/2\right),\hskip 3.55658pt\mu\mid\sigma^{2}\sim{\mathcal{N}}\left(\mu_{0},\sigma^{2}/\kappa_{0}\right)$ (2)
for analyst-specified hyperparameters $(\nu_{0},\kappa_{0},\mu_{0},\sigma_{0}^{2}).$ Here, IG denotes the inverse-gamma distribution. The prior distribution in (2) is conjugate for the normal model, which facilitates computation. Using (2), the posterior distribution $p(\mu,\sigma^{2}|\bar{Y}^{*},S^{2*})$ can be estimated using Markov chain Monte Carlo techniques.
Bernstein & Sheldon (2018, 2019) present Gibbs samplers for similar inferential tasks, in which one samples from the full conditional of each unknown parameter in $(\mu,\sigma^{2},\bar{Y},S^{2})$ at each iteration. However, these samplers rely on an approximation to the full conditional distribution of $S^{2}$, namely that it is normally distributed. Clearly this is not the case, since $S^{2}$ must be positive. Thus, we come to our first innovation: we derive a Gibbs sampler that utilizes the exact full conditional distribution of $S^{2}$, thereby eschewing the normal approximation for $S^{2}$. Specifically, when $\varepsilon_{2}<2$, the full conditional for $S^{2}$ is a distribution we call a truncated gamma mixture (TGM). [^1] For the distribution of $S^{2}$ to be a TGM, we require $\sigma^{2}<(n-1)/(2n\varepsilon_{2})$. We show in Corollary 1 that in this setting we can guarantee $\sigma^{2}\leq 1/4$, which implies $\sigma^{2}<(n-1)/(2n\varepsilon_{2})$ automatically when $\varepsilon_{2}<2$. Within the Gibbs sampler, we reject and resample any draws for which $\sigma^{2}\geq(n-1)/(2n\varepsilon_{2})$.
@@
$X\sim\textsf{TGM}(\alpha,\beta,\lambda,\tau)$ for $\alpha>0$, $\tau\in{\mathbb{R}}$, and $\beta>\lambda\geq 0$ if its distribution is as follows. When $\tau\leq 0$, $X\sim\textsf{Gamma}(\alpha,\beta+\lambda)$ and when when $\tau>0$, it has probability density function
$p(x)=\begin{cases}\pi_{1}\frac{(\beta-\lambda)^{\alpha}}{\gamma(\alpha,(\beta-\lambda)\tau)}x^{\alpha-1}e^{-(\beta-\lambda)x},&\mbox{if }x\leq\tau;\\ \pi_{2}\frac{(\beta+\lambda)^{\alpha}}{\Gamma(\alpha,(\beta+\lambda)\tau)}x^{\alpha-1}e^{-(\beta+\lambda)x}&\mbox{if }x>\tau,\end{cases}$ (3)
@@
$\begin{matrix}\pi_{1}&=\frac{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}}{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}+e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}},\\ \pi_{2}&=\frac{e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}}{e^{-\lambda\tau}\frac{\gamma(\alpha,(\beta-\lambda)\tau)}{(\beta-\lambda)^{\alpha}}+e^{\lambda\tau}\frac{\Gamma(\alpha,(\beta+\lambda)\tau)}{(\beta+\lambda)^{\alpha}}}.\end{matrix}$ (4)
@@
Here, $\gamma(\alpha,x)$ and $\Gamma(\alpha,x)$ are the lower and upper incomplete gamma functions. We provide an algorithm for sampling from the TGM in Appendix B.
@@
Derivations for all full conditionals for the Gibbs sampler are provided in Appendix C. The full conditionals for $\mu$ and $\sigma^{2}$ are identical to those of the nonprivate setting. To obtain the full conditional for $\bar{Y}$, we follow the strategy of Bernstein & Sheldon (2018) based on the Bayesian LASSO (Park & Casella, 2008) and use the fact that $\bar{Y}^{*}\sim\textsf{Lap}(\bar{Y},1/(\varepsilon_{1}n))$ is equal in distribution to $\bar{Y}^{*}\sim{\mathcal{N}}(\bar{Y},\omega^{2})$ for $\omega^{2}\sim\textsf{Exp}(\varepsilon_{1}^{2}n^{2}/2)$. The full conditional for $\bar{Y}^{*}$ is Gaussian, the full conditional for $1/\omega^{2}$ is inverse-Gaussian, and the full conditional for $S^{2}$ is a TGM. The parameters of the full conditionals and an explicit algorithm for our Gibbs sampler are provided in Appendix D.
@@
A major advantage of this sampler is its runtime. The computational complexity of this sampler is $\mathcal{O}(T)$, where $T$ is the number of Gibbs iterations. The runtime does not scale with $n$. The Gibbs sampler of Bernstein & Sheldon (2018) applied to the univariate Gaussian setting also has computational complexity $\mathcal{O}(T)$. This sampler, however, requires a Gaussian approximation for $S^{2}$, which, particularly when $\sigma^{2}$ is near zero or $n$ is small, can lead to unreliable results. The sampler of Ju et al. (2022) does not require any approximations; however, it has computational complexity $\mathcal{O}(nT)$, making it potentially infeasible for large $n$. Our proposed sampler achieves both the exact sampling of Ju et al. (2022) and the $\mathcal{O}(T)$ complexity of Bernstein & Sheldon (2018).
3 ENFORCING CONSTRAINTS